# Future-Back Threat Modeling: A Foresight-Driven Security Framework

### v1

@@ Abstract
As first articulated in [isc22025futureback] and later expanded through polycritical foresight discussions [realitystudies2025polycritical], the Future-Back approach integrates epistemic foresight into cyber threat modeling, aligning with ISACA’s revisited perspective on systemic assurance [isaca2025threatmodelingwhitepaper]. Collier (2020) further demonstrated how backcasting can reveal systemic weaknesses within cyber threat intelligence ecosystems, supporting its epistemic role in foresight-driven security [collier2020backcasting].
@@ 1.2 Theoretical Gap
Strategic foresight and threatcasting methods, especially vanatta2019threatcasting and vescent2018shifting, offer narrative and variable-based tools to anticipate long-term change. However, they usually end before providing actionable engineering artifacts.
@@ 2 Related Work – Where FBTM Extends Existing Thinking
In contrast, approaches such as threatcasting [vanatta2019threatcasting] As further conceptual grounding, Vescent and Blakley (2018) explored the role of strategic foresight in evolving security paradigms, providing an early academic basis for integrating foresight into cybersecurity planning [vescent2018shifting].
and strategic foresight [vescent2018shifting] explore possible futures, but stop short of turning those futures into testable security controls.
@@ 4.1 Phase I – Foresight and Assumption Mapping
This phase inverts the direction of traditional threat modeling by beginning with future states rather than current systems. The objective is to articulate the beliefs, expectations, and biases that shape an organization’s sense of safety — then make them explicit and testable.
@@ 4.1.2 Step 2 – Future-State Assumption Mapping
This integrated map combines global foresight—EU-level systemic threats identified by ENISA [enisa2023foresight]—with domain-specific foresight derived from the firm’s projected 2040 growth trajectory. Each assumption represents a testable belief about the organization’s future safety conditions. Fragility ratings reveal the intersection between global uncertainty and enterprise-level stressors, indicating where systemic volatility may amplify organizational vulnerabilities.
@@ 6.1 Step 5a – Governance and Legal Gate
After each epistemic test, results are reviewed through a governance gate that mirrors the Assess–Authorize–Monitor phases of the NIST Risk Management Framework[nist_sp80037r2]. . This isn’t a bureaucratic stop; it’s where discovery becomes accountability.
@@
    Policy connection — mapping what we learned back to existing control frameworks (NIST SP 800-53, ISO 27001, GDPR) [nist2020sp80053, iso27001_2022, gdpr2016].
@@ 6.3 Step 6 – Operational Response and Learning
Once governance decisions are made, the cycle loops back into operations. Here, FBTM connects directly with NIST SP 800-61r2 Incident Response [nist2012sp80061r2]—but treats every incident not just as an event to contain, but as an epistemic experiment to learn from.
@@ 7 Phase IV – Synthesis and Strategic Implications
In this sense, FBTM does not stand outside of existing standards like NIST CSF or ISO 27001 [nistcsf2024, iso27001_2022]. Rather, it acts as an epistemic engine inside them, converting static controls into dynamic knowledge.

### v2

@@ Abstract
As first articulated in ISC2 Insight Future-Back Threat Modeling [isc22025futureback], the Future-Back approach integrates epistemic foresight into cyber threat modeling, aligning with ISACA’s revisited perspective on systemic assurance [isaca2025threatmodelingwhitepaper]. Collier (2020) further demonstrated how backcasting can reveal systemic weaknesses within cyber threat intelligence ecosystems, supporting its epistemic role in foresight-driven security [collier2020backcasting].
@@ 1.2 Theoretical Gap
Strategic foresight methods—such as those built on scenario exploration and backcasting—provide narrative and variable-driven approaches to anticipating long-term change. Yet they rarely yield the kind of engineering-ready artifacts or assumption-testing mechanisms required for operational cybersecurity practice.
@@ 2 Related Work – Where FBTM Extends Existing Thinking
In contrast, many foresight-oriented approaches focus on constructing plausible future scenarios but seldom examine whether the assumptions underlying those scenarios would hold under adversarial or unexpected conditions. They help organizations imagine what might happen, but they do not translate those futures into testable security controls or mechanisms for validating belief structures.
@@ 4.1 Phase I – Foresight and Assumption Mapping
This phase inverts the direction of traditional threat modeling by beginning with future states rather than current systems. The objective is to articulate the beliefs, expectations, and biases that shape an organization’s sense of safety — then make them explicit and testable. This reflects Karl Popper’s view that progress in knowledge arises from the falsification of assumptions rather than their confirmation [popper1959falsification], and aligns with John Robinson’s insight that backcasting reveals how future scenarios depend on present-state assumptions [robinson1990backcasting].
@@ 4.1.2 Step 2 – Future-State Assumption Mapping
This integrated map combines global foresight—EU-level systemic threats identified by ENISA [enisa2030foresight]—with domain-specific foresight derived from the firm’s projected 2040 growth trajectory. Each assumption represents a testable belief about the organization’s future safety conditions. Fragility ratings reveal the intersection between global uncertainty and enterprise-level stressors, indicating where systemic volatility may amplify organizational vulnerabilities.
@@ 6.1 Step 5a – Governance and Legal Gate
After each epistemic test, results are reviewed through a governance gate that mirrors the Assess–Authorize–Monitor phases of the NIST Risk Management Framework. . This isn’t a bureaucratic stop; it’s where discovery becomes accountability.
@@
    Policy connection — mapping what we learned back to existing control frameworks (NIST SP 800-53, ISO 27001,) [nist2020sp80053].
@@ 6.3 Step 6 – Operational Response and Learning
Once governance decisions are made, the cycle loops back into operations. Here, FBTM connects directly with NIST SP 800-61r2 Incident Response —but treats every incident not just as an event to contain, but as an epistemic experiment to learn from.
@@ 7 Phase IV – Synthesis and Strategic Implications
In this sense, FBTM does not stand outside of existing standards like NIST CSF or ISO 27001 [nistcsf2024]. Rather, it acts as an epistemic engine inside them, converting static controls into dynamic knowledge.